# A consensus-constrained parsimonious Gaussian mixture model for clustering hyperspectral images

### v1

@@ Abstract
The use of hyperspectral imaging to investigate food samples has grown due to the improved performance and lower cost of spectroscopy instrumentation. Food engineers use hyperspectral images to classify the type and quality of a food sample, typically using classification methods. In order to train these methods, every pixel in each training image needs to be labelled. Typically, computationally cheap threshold-based approaches are used to label the pixels, and classification methods are trained based on those labels. However, threshold-based approaches are subjective and cannot be generalized across hyperspectral images taken in different conditions and of different foods. Here a consensus-constrained parsimonious Gaussian mixture model (ccPGMM) is proposed to label pixels in hyperspectral images using a model-based clustering approach. The ccPGMM utilizes available information on the labels of a small number of pixels and the relationship between those pixels and neighbouring pixels as constraints when clustering the rest of the pixels in the image. A latent variable model is used to represent the high-dimensional data in terms of a small number of underlying latent factors. To ensure computational feasibility, a consensus clustering approach is employed, where the data are divided into multiple randomly selected subsets of variables and constrained clustering is applied to each data subset; the clustering results are then consolidated across all data subsets to provide a consensus clustering solution. The ccPGMM approach is applied to simulated datasets and real hyperspectral images of three types of puffed cereal, corn, rice, and wheat. Improved clustering performance and computational efficiency are demonstrated when compared to other current state-of-the-art approaches.
@@ 1 Introduction
A major challenge in labelling the pixels of the hyperspectral images is that each image has a large number of pixels and for each pixel, a large amount of reflectance information is captured in a continuous $p$-dimensional spectrum, whose wavelengths are highly correlated. Further, often threshold and $k$-means clustering approaches are applied to each hyperspectral image individually, which is impractical as the number of images increases and does not allow for the borrowing of information across images. Also, the threshold approach and $k$-means clustering with 2 clusters work well only when there is exactly one type of food sample in a hyperspectral image. As the number of food samples in a given image increases, multiple thresholds need to be selected from the histogram, which is even more challenging. Overall, threshold based approaches are subjective and are not easily generalisable across hyperspectral images.
@@
Here, we propose a novel model-based clustering approach, a consensus-constrained parsimonious Gaussian mixture model (ccPGMM), to label the pixels of multiple hyperspectral images simultaneously in a computationally feasible manner while accounting for uncertainty. A key aspect of ccPGMM is that it involves a constrained parsimonious Gaussian mixture model (Melnykov et al., , 2016) in which information regarding the types of food samples present in the hyperspectral images along with information about some pixels is incorporated as constraints to perform informed clustering. Further, as hyperspectral images have a large number of pixels, each of which has an associated high-dimensional spectrum of $p$ strongly correlated variables, ccPGMM employs a parsimonious Gaussian mixture model (PGMM) which describes the high-dimensional data using a small number of latent variables. To enable computationally efficient implementation, ccPGMM adopts a consensus approach, whereby the high-dimensional spectrum is divided into randomly selected subsets of $d<p$ variables, and a constrained PGMM is fitted to each subset of variables on all pixels. The clustering solutions of the constrained PGMMs fitted to each subset are consolidated to provide a final consensus clustering allocation for each pixel. The performance of the proposed ccPGMM approach is demonstrated, and compared to other state-of-the-art approaches, through thorough simulation studies and application to real hyperspectral images of puffed cereals (Gowen et al., , 2019).
In what follows, Section 2 gives details of the motivating dataset containing hyperspectral images of puffed cereals. Section 3 outlines the ccPGMM approach and discusses its inference and implementation details. Section 4 delineates thorough simulation studies that explore the performance of ccPGMM under different settings. Section 5 discusses the application of ccPGMM to the motivating puffed cereal images and Section 6 summarises the contributions of ccPGMM and gives an overview of future research directions. The R (R Core Team, , 2023) code to implement ccPGMM is available at this repository GitHub.
@@ 2 Hyperspectral images of puffed cereals
The ccPGMM approach is motivated by the need to simultaneously label pixels in nine hyperspectral images of three types of puffed cereal, wheat, corn, and rice (Gowen et al., , 2019). Each pixel in an image records the near-infrared (NIR) spectrum, with $p=101$ equally spaced variables spanning the wavelengths in the 880-1720nm interval. Each image $l=1,\ldots,L=9$ contains only one puffed cereal and is a three-dimensional tensor of dimension $S_{l}\times T_{l}\times p$ where $S_{l}$ and $T_{l}$ are the numbers of row and column pixels respectively, which vary across the images as they are of different dimensions (details are given in Appendix 8). Figure 1 shows a greyscale image of one hyperspectral image of each puffed cereal type, where, for each pixel, the intensity of the color is given by the average of the NIR spectrum.
For the purpose of ccPGMM, the pixels are assumed to be independent and so the three-dimensional tensor of each image is represented by a rectangular dataset of dimension $(S_{l}T_{l})\times p$. These $L=9$ rectangular datasets are then collated to form a single dataset of dimension $N\times p$, where $N=\sum_{l=1}^{L}S_{l}T_{l}$. Here, the final dataset of nine images results in $N=28039$ pixels, each measured over $p=101$ wavelengths which are highly correlated and contain spectral noise.
The objective is to simultaneously label all pixels in all nine images as either background or a cereal type, taking into account available information on the images (i.e., it is known which cereal type is in each image) and on some pixels (e.g., corner pixels are background), in a computationally feasible manner.
@@
Figure 1: Greyscale images of one hyperspectral image of each puffed cereal type. For each pixel, the intensity of the color corresponds to the average of its NIR spectrum.
@@ 3.1 Parsimonious Gaussian mixture models (PGMM)
Factor analysis is a latent variable model that represents a large number $p$ of correlated variables using a smaller number $q\ll p$ of underlying latent factors. In a single factor analysis model (McLachlan and Peel, , 2000), each observation $\mathbf{x}_{n}\in\mathbb{R}^{p}$ for $n=1,\ldots,N$ in data $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ of dimension $N\times p$ is modelled as $\mathbf{x}_{n}=\bm{\mu}+\bm{\Lambda}\mathbf{u}_{n}+\bm{\varepsilon}_{n}$ where $\bm{\mu}$ is the mean, $\bm{\Lambda}$ is the $p\times q$ loadings matrix and $\mathbf{u}_{n}\sim\mathcal{N}_{q}(\mathbf{0},\mathbf{I})$ and $\bm{\varepsilon}_{n}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{\Psi})$ are the latent factor score and specific factor, respectively, for observation $n$. Therefore, $\mathbf{x}_{n}\sim\mathcal{N}_{p}(\bm{\mu},\bm{\Sigma})$, where ${\bm{\Sigma}}=\bm{\Lambda\Lambda^{\top}+\Psi}$ provides a parsimonious covariance structure.
The parsimonious Gaussian mixture model (PGMM) (McLachlan and Peel, , 2004; McNicholas and Murphy, , 2008; McNicholas et al., , 2010) is a finite mixture of $G$ factor analysis models such that each mixture component $g$ has a multivariate Gaussian distribution with mean $\bm{\mu}_{g}$ and covariance $\bm{\Lambda}_{g}\bm{\Lambda}_{g}^{\top}+\bm{\Psi}_{g}$. Denoting the probability of membership of mixture component $g$ as $\tau_{g}$, under PGMM, we have ${f}(\mathbf{x}_{n})=\sum_{g=1}^{G}\tau_{g}\>\mathcal{N}_{p}(\mathbf{x}_{n};\bm{\mu}_{g},\bm{\Lambda}_{g},\bm{\Psi}_{g}).$ By constraining $\bm{\Lambda}_{g}$ and $\bm{\Psi}_{g}$ to be the same or different across components, a family of PGMMs have been developed (McNicholas and Murphy, , 2008) with inference via an alternating expectation-conditional maximization(AECM, Ghahramani et al., , 1996; McLachlan et al., , 2003). These PGMMs facilitate model-based clustering of high-dimensional data.
@@ 3.2 Constrained parsimonious Gaussian mixture models
When labelling pixels in a set of hyperspectral images, there is often information about the type of food sample(s) present in some of the hyperspectral images, along with indirect information on the labels of some pixels in the images. For example, in Figure 2, four blocks of pixels are highlighted in three different colours in the two images. The pixels in the blue blocks located at the corner of the images are assumed to belong to the background. On the other hand, the pixels in the yellow and green blocks are located in the center of their images, where the corresponding cereal is observed to be situated. Pixels in the blue blocks should be clustered together, as must the pixels in the yellow block and the pixels in the green block; this information imposes what are known as positive constraints. Also, the pixels in the blue blocks should not be placed in the same cluster as pixels in the yellow or green blocks, and vice versa; this information imposes what are known as negative constraints.
Melnykov et al., (2016) proposed a constrained Gaussian mixture model that uses such additional information about some observations as constraints to perform informed model-based clustering. However, the high-dimensionality of the hyperspectral imaging data poses computational challenges for this approach. Here, we build on Melnykov et al., (2016) and allow for constraints to be incorporated when clustering using a PGMM. The resulting constrained parsimonious Gaussian mixture model (constrained-PGMM) facilitates clustering and therefore labelling of the pixels of the hyperspectral images.
@@
Figure 2: An illustration of positive ($+$) and negative (--) constraints. Four blocks of three types of pixels are highlighted. Pixels in the blue blocks (assumed to be the background) must be clustered together, as must the pixels in the yellow block (wheat) and the pixels in the green block (corn). However, pixels in the blue blocks must not be clustered with those in the yellow block or with those in the green block, and vice versa.
Under the constrained-PGMM, the observed data log-likelihood is
$\ell_{o}(\mathbf{X})=\sum_{n=1}^{N}\log\sum_{g=1}^{G}\tau_{g}\>\mathcal{N}_{p}(\mathbf{x}_{n};\bm{\mu}_{g},\bm{\Lambda}_{g},\bm{\Psi}_{g}).$ (1)
@@
As (1) is difficult to maximize directly, an AECM algorithm is used to fit the constrained-PGMM. This requires the introduction of a latent component indicator $\mathbf{z}_{n}=(z_{n1},\ldots,z_{nG})^{\prime}$, for $n=1,\ldots,N$, where $z_{ng}=1$ if observation $n$ belongs to cluster $g$ or 0 otherwise. The AECM algorithm works in two cycles and allows for a different definition of the complete data in each cycle. In the first cycle, the component indicators $\mathbf{Z}$ = $\mathbf{z}_{1}$,…, $\mathbf{z}_{N}$} are assumed to be the missing data. The complete data log-likelihood is then
$\ell_{c}(\mathbf{X},\mathbf{Z})=\sum_{n=1}^{N}\sum_{g=1}^{G}{z}_{ng}[\log\tau_{g}+\log\>\mathcal{N}_{p}(\mathbf{x}_{n};\bm{\mu}_{g},\bm{\Lambda}_{g},\bm{\Psi}_{g})].$ (2)
@@
In the E-step of the first cycle, the available constraints are incorporated into PGMM. Let $B_{1}$ be the set of $J$ pixels (indexed by $j$) in the blue blocks shown in Figure 2 which are known to be background and must be clustered together. Let $B_{2}$ be the set of $O$ pixels (indexed by $o$) in the yellow block of Figure 2 which are known to be wheat and should be clustered together. These are positive constraints (denoted $+$). However, the pixels in $B_{1}$ should not be clustered together with the pixels in $B_{2}$. This is a negative constraint (denoted --). Thus the posterior probability $\hat{{z}}_{{B_{1}},g}^{+\atop-}$ of the pixels in $B_{1}$ belonging to cluster $g$, given the positive and negative constraints $+\atop-$ is
@@
$\centering\indent\indent\hat{{z}}_{{B_{1}},g}^{+\atop-}=\frac{\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{g}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{g},\hat{\bm{\Lambda}}_{g}\>\hat{\bm{\Lambda}}_{g}^{\top}\>+\>\hat{\bm{\Psi}}_{g})\sum\limits_{\begin{subarray}{c}f=1\\ f\neq g\end{subarray}}^{G}\prod\limits_{\begin{subarray}{c}o\in{B_{2}}\\ o=1\end{subarray}}^{O}\hat{\tau}_{f}\mathcal{N}_{p}(\mathbf{x}_{o};\hat{\bm{\mu}}_{f},\hat{\bm{\Lambda}}_{f}\>\hat{\bm{\Lambda}}_{f}^{\top}\>+\>\hat{\bm{\Psi}}_{f})}{\displaystyle\sum_{g^{{}^{\prime}}=1}^{G}\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{g^{{}^{\prime}}}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{g^{{}^{\prime}}},\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}\>\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}^{\top}\>+\>\hat{\bm{\Psi}}_{g^{{}^{\prime}}})\sum\limits_{\begin{subarray}{c}f=1\\ f\neq g^{{}^{\prime}}\end{subarray}}^{G}\prod\limits_{\begin{subarray}{c}o\in{B_{2}}\\ o=1\end{subarray}}^{O}\hat{\tau}_{f}\mathcal{N}_{p}(\mathbf{x}_{o};\hat{\bm{\mu}}_{f},\hat{\bm{\Lambda}}_{f}\>\hat{\bm{\Lambda}}_{f}^{\top}\>+\>\hat{\bm{\Psi}}_{f})},\@add@centering$ (3)

### v2

@@ Abstract
The use of hyperspectral imaging to investigate food samples has grown due to the improved performance and lower cost of instrumentation. Food engineers use hyperspectral images to classify the type and quality of a food sample, typically using classification methods. In order to train these methods, every pixel in each training image needs to be labelled. Typically, computationally cheap threshold-based approaches are used to label the pixels, and classification methods are trained based on those labels. However, threshold-based approaches are subjective and cannot be generalized across hyperspectral images taken in different conditions and of different foods. Here a consensus-constrained parsimonious Gaussian mixture model (ccPGMM) is proposed to label pixels in hyperspectral images using a model-based clustering approach. The ccPGMM utilizes information that is available on some pixels and specifies constraints on those pixels belonging to the same or different clusters while clustering the rest of the pixels in the image. A latent variable model is used to represent the high-dimensional data in terms of a small number of underlying latent factors. To ensure computational feasibility, a consensus clustering approach is employed, where the data are divided into multiple randomly selected subsets of variables and constrained clustering is applied to each data subset; the clustering results are then consolidated across all data subsets to provide a consensus clustering solution. The ccPGMM approach is applied to simulated datasets and real hyperspectral images of three types of puffed cereal, corn, rice, and wheat. Improved clustering performance and computational efficiency are demonstrated when compared to other current state-of-the-art approaches.
@@ 1 Introduction
A major challenge in labelling the pixels of the hyperspectral images is that each image has a large number of pixels and for each pixel, a large amount of highly correlated reflectance information is captured in a continuous $p$-dimensional spectrum. Further, often threshold and $k$-means clustering approaches are applied to each hyperspectral image individually, which is impractical as the number of images increases and does not allow for the borrowing of information across images. Also, the threshold approach and $k$-means clustering with 2 clusters often only work well when there is exactly one type of food sample in a hyperspectral image. As the number of food samples in a given image increases, multiple thresholds need to be selected from the histogram, which is even more challenging. Overall, threshold-based approaches are subjective and are not easily generalisable across hyperspectral images.
@@
Here, we propose a novel model-based clustering approach, a consensus-constrained parsimonious Gaussian mixture model (ccPGMM), to label the pixels of multiple hyperspectral images simultaneously in a computationally feasible manner while accounting for uncertainty. A key aspect of ccPGMM is that it involves a constrained parsimonious Gaussian mixture model (constrained-PGMM) in which information regarding the types of food samples present in the hyperspectral images along with information about some pixels is incorporated as constraints to perform informed clustering. Further, as hyperspectral images have a large number of pixels, each of which has an associated high-dimensional spectrum of $p$ strongly correlated variables, ccPGMM employs a parsimonious Gaussian mixture model (PGMM) which describes the high-dimensional data using a small number of latent variables. To enable computationally efficient implementation, ccPGMM adopts a consensus approach, whereby the high-dimensional spectrum is divided into randomly selected subsets of $d<p$ variables, and a constrained-PGMM is fitted to each subset of variables on all pixels. The clustering solutions of the constrained-PGMM fitted to each subset are consolidated to provide a final consensus clustering allocation for each pixel. The performance of the proposed ccPGMM approach is demonstrated, and compared to other state-of-the-art approaches, through thorough simulation studies and application to real hyperspectral images of puffed cereals (Gowen et al., , 2019).
In what follows, Section 2 gives details of the motivating dataset containing hyperspectral images of puffed cereals. Section 3 outlines the ccPGMM approach and discusses its inference and implementation details. Section 4 delineates thorough simulation studies that explore the performance of ccPGMM under different settings. Section 5 discusses the application of ccPGMM to the motivating puffed cereal images and Section 6 summarises the contributions of ccPGMM and gives an overview of future research directions. The R (R Core Team, , 2023) code to implement ccPGMM is available at github.com/GaneshBabu1604/ccPGMM.git.
@@ 2 Hyperspectral images of puffed cereals
The ccPGMM approach is motivated by the need to simultaneously label pixels in nine hyperspectral images of three types of puffed cereal: wheat, corn, and rice (Gowen et al., , 2019). Each pixel in an image records the near-infrared (NIR) spectrum, with $p=101$ equally spaced variables spanning the wavelengths in the 880-1720nm interval. Each image $l=1,\ldots,L$, where $L=9$ contains only one puffed cereal and is a three-dimensional tensor of dimension $S_{l}\times T_{l}\times p$ where $S_{l}$ and $T_{l}$ are the numbers of row and column pixels respectively, which vary across the images as they are of different dimensions (details are given in Appendix A). Figure 1 shows a greyscale image of one hyperspectral image of each puffed cereal type, where, for each pixel, the intensity of the colour is given by the average of the reflectance information captured in the NIR spectrum.
For the purpose of ccPGMM, the pixels are assumed to be independent and so the three-dimensional tensor of each image $l$ is represented by a rectangular dataset of dimension $(S_{l}T_{l})\times p$. These $L=9$ rectangular datasets are then collated to form a single dataset of dimension $N\times p$, where $N=\sum_{l=1}^{L}S_{l}T_{l}$. Here, the final dataset of nine images results in $N=28039$ pixels, each measured over $p=101$ wavelengths which are highly correlated and contain spectral noise.
The objective is to simultaneously label all pixels in all nine images as either background or a cereal type, taking into account available information on the images (i.e., it is known which cereal type is in each image) and on some pixels (e.g., some corner pixels are likely to be background and should be clustered together), in a computationally feasible manner. To demonstrate the utility of the proposed ccPGMM approach, the method is used in a supervised manner to classify the pixels of an additional hyperspectral image that contains multiple puffed cereal grains of potentially different type. Details of the dimensions of this multi-grain image are given in Appendix A.
@@
Figure 1: Greyscale images of one hyperspectral image of each puffed cereal type. For each pixel, the intensity of the color corresponds to the average of the reflectance information captured in the NIR spectrum.
@@ 3.1 Parsimonious Gaussian mixture models (PGMM)
Factor analysis is a latent variable model that represents a large number $p$ of correlated variables using a smaller number $q\ll p$ of underlying latent factors. In a single factor analysis model (McLachlan and Peel, , 2000), each observation $\mathbf{x}_{n}\in\mathbb{R}^{p}$ for $n=1,\ldots,N$ in data $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ of dimension $N\times p$ is modelled as $\mathbf{x}_{n}=\bm{\mu}+\bm{\Lambda}\mathbf{u}_{n}+\bm{\varepsilon}_{n}$, where $\bm{\mu}$ is the mean, $\bm{\Lambda}$ is the $p\times q$ loadings matrix and $\mathbf{u}_{n}\sim\mathcal{N}_{q}(\mathbf{0},\mathbf{I})$ and $\bm{\varepsilon}_{n}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{\Psi})$ are the latent factor score and specific factor, respectively, for observation $n$. Therefore, $\mathbf{x}_{n}\sim\mathcal{N}_{p}(\bm{\mu},\bm{\Sigma})$, where ${\bm{\Sigma}}=\bm{\Lambda\Lambda^{\top}+\Psi}$ provides a parsimonious covariance structure.
The parsimonious Gaussian mixture model (PGMM) (McLachlan and Peel, , 2004; McNicholas and Murphy, , 2008; McNicholas et al., , 2010) is a finite mixture of $G$ factor analysis models such that each mixture component $g$ has a multivariate Gaussian distribution with mean $\bm{\mu}_{g}$ and covariance $\bm{\Sigma}_{g}=\bm{\Lambda}_{g}\bm{\Lambda}_{g}^{\top}+\bm{\Psi}_{g}$. Denoting the probability of membership of mixture component $g$ as $\tau_{g}$, under PGMM, we have ${f}(\mathbf{x}_{n})=\sum_{g=1}^{G}\tau_{g}\>\mathcal{N}_{p}(\mathbf{x}_{n};\bm{\mu}_{g},\bm{\Lambda}_{g},\bm{\Psi}_{g}).$ By constraining $\bm{\Lambda}_{g}$ and $\bm{\Psi}_{g}$ to be the same or different across components, a family of PGMMs has been developed (McNicholas and Murphy, , 2008) with inference via an alternating expectation-conditional maximization(AECM, Ghahramani et al., , 1996; McLachlan et al., , 2003). These PGMMs facilitate model-based clustering of high-dimensional data.
@@ 3.2 Constrained parsimonious Gaussian mixture models
When labelling pixels in a set of hyperspectral images, there is often information about the type of food sample(s) present in some of the hyperspectral images, along with indirect information on the likely labels of some pixels in the images. For example, in Figure 2, four blocks of pixels are highlighted in three different colours in the two images. The pixels in the blue blocks located at the corner of the images are assumed to belong to the same cluster, corresponding to the background. On the other hand, the pixels in the yellow and green blocks are located in the center of their images, where the corresponding cereal is observed to be situated. Pixels in the blue blocks should be clustered together, as must the pixels in the yellow block and the pixels in the green block; this information imposes what are known as positive constraints. Also, the pixels in the blue blocks should not be placed in the same cluster as pixels in the yellow or green blocks, and vice versa; this information imposes what are known as negative constraints. It is not necessary to specify constraints in all images if the user prefers or is uncertain, e.g., as may occur in a case where multiple puffed cereal grains of potentially different types are present in a single training image. As highlighted by Melnykov et al., (2016), the positive and negative constraints relate to pixels belonging to the same or different clusters, rather than directly specifying pixel labels; unless in a trivial case, assigning labels to pixels prior to model fitting is questionable since the components are yet to be formed and there is therefore inherent uncertainty.
Melnykov et al., (2016) proposed a constrained Gaussian mixture model that uses such additional information about some observations as constraints to perform informed model-based clustering. However, the high-dimensionality of the hyperspectral imaging data poses computational challenges for this approach. Here, we build on Melnykov et al., (2016) and allow for constraints to be incorporated when clustering using a PGMM. The resulting constrained parsimonious Gaussian mixture model (constrained-PGMM) facilitates clustering and therefore labelling of the pixels of the hyperspectral images. Denoting the parameter vector $\bm{\Theta}=\{\tau_{1},\ldots,\tau_{G},\bm{\mu}_{1},\ldots,\bm{\mu}_{G},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{G},\bm{\Psi}_{1},\ldots,\bm{\Psi}_{G}\}$, under the constrained-PGMM, the observed data log-likelihood is
@@
Figure 2: An illustration of positive ($+$) and negative (--) constraints. Four blocks of three types of pixels are highlighted. Pixels in the blue blocks must be clustered together, as must the pixels in the yellow block and the pixels in the green block. However, pixels in the blue blocks must not be clustered with those in the yellow block or with those in the green block, and vice versa.
$\ell_{o}(\bm{\Theta};\mathbf{X})=\sum_{n=1}^{N}\log\sum_{g=1}^{G}\tau_{g}\>\mathcal{N}_{p}(\mathbf{x}_{n};\bm{\mu}_{g},\bm{\Lambda}_{g},\bm{\Psi}_{g}).$ (1)
@@
As (1) is difficult to maximize directly, an AECM algorithm is used to fit the constrained-PGMM. This requires the introduction of a latent component indicator $\mathbf{z}_{n}=(z_{n1},\ldots,z_{nG})^{\prime}$, for $n=1,\ldots,N$, where $z_{ng}=1$ if observation $n$ belongs to cluster $g$ or 0 otherwise. The AECM algorithm works in two cycles and allows for a different definition of the complete data in each cycle. In the first cycle, the component indicators $\mathbf{Z}$ = $\mathbf{z}_{1}$,…, $\mathbf{z}_{N}$} are assumed to be the missing data. Further, $\bm{\Theta}$ is partitioned as $(\bm{\Theta}_{1},\bm{\Theta}_{2})$ where $\bm{\Theta}_{1}$ denotes the mixing proportions $\tau_{g}$ and the mean parameters $\bm{\mu}_{g}$ ($g=1,\ldots,G$), and $\bm{\Theta}_{2}$ denotes the loadings $\bm{\Lambda}_{g}$ and variances $\bm{\Psi}_{g}$ for $g=1,\ldots,G$. The complete data log-likelihood in the first cycle is then
$\ell_{c}(\bm{\Theta};\mathbf{X},\mathbf{Z})=\sum_{n=1}^{N}\sum_{g=1}^{G}{z}_{ng}[\log\tau_{g}+\log\>\mathcal{N}_{p}(\mathbf{x}_{n};\bm{\mu}_{g},\bm{\Lambda}_{g},\bm{\Psi}_{g})]$ (2)
and the expected complete data log-likelihood is
$Q_{1}(\bm{\Theta})\!\!$ $=$ $\!\!\sum_{n=1}^{N}\sum_{g=1}^{G}E[{z}_{ng};\mathbf{x}_{n}]\Bigg{[}\log\tau_{g}-\frac{p}{2}\log(2\pi)-\frac{1}{2}\log(|\bm{\Lambda}_{g}\>\bm{\Lambda}_{g}^{\top}\>+\>\bm{\Psi}_{g}|)$
$-\frac{1}{2}[(\mathbf{x}_{n}-\bm{\mu}_{g})^{\top}(\bm{\Lambda}_{g}\>\bm{\Lambda}_{g}^{\top}\>+\>\bm{\Psi}_{g})^{-1}(\mathbf{x}_{n}-\bm{\mu}_{g})]\Bigg{]}.$
The posterior probability ${\hat{z}_{ng}}$ of observation $n$ belonging to cluster $g$ is
$E[{z}_{ng};\mathbf{x}_{n}]$ $=$ $\hat{{z}}_{ng}\>=\>\Pr({z}_{ng}=1;\mathbf{x}_{n},\hat{\tau}_{g},\hat{\bm{\mu}}_{g},\hat{\bm{\Lambda}}_{g},\hat{\bm{\Psi}}_{g})=\frac{\hat{\tau}_{g}\>\mathcal{N}_{p}(\mathbf{x}_{n};\hat{\bm{\mu}}_{g},\hat{\bm{\Lambda}}_{g}\>\hat{\bm{\Lambda}}_{g}^{\top}\>+\>\hat{\bm{\Psi}}_{g})}{\sum_{g^{{}^{\prime}}=1}^{G}\hat{\tau}_{g^{{}^{\prime}}}\>\mathcal{N}_{p}(\mathbf{x}_{n};\hat{\bm{\mu}}_{g^{{}^{\prime}}},\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}\>\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}^{\top}\>+\>\hat{\bm{\Psi}}_{g^{{}^{\prime}}})}$ (3)
@@
where $\hat{\cdot}$ denotes an initial or current parameter estimate, as relevant. Let $B_{1}$ be the set of $J$ pixels (indexed by $j$) which must be clustered together, shown in the blue blocks in Figure 2. This is a positive constraint (denoted $+$). The posterior probability $\hat{z}_{{B_{1}}g}^{+}$ of the pixels in $B_{1}$ belonging to cluster $g$ given the positive constraints ($+$) is
@@
$\centering\indent\indent\hat{{z}}_{{B_{1}},g}^{+}=\frac{\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{g}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{g},\hat{\bm{\Lambda}}_{g}\>\hat{\bm{\Lambda}}_{g}^{\top}\>+\>\hat{\bm{\Psi}}_{g})}{\displaystyle\sum_{g^{{}^{\prime}}=1}^{G}\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{g^{{}^{\prime}}}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{g^{{}^{\prime}}},\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}\>\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}^{\top}\>+\>\hat{\bm{\Psi}}_{g^{{}^{\prime}}})}.\@add@centering$
Let $B_{2}$ be the set of $K$ pixels (indexed by $k$) which must be clustered together, shown in the yellow block of Figure 2; these are positive constraints (denoted $+$). However, the pixels in $B_{1}$ should not be clustered together with the pixels in $B_{2}$; this is a negative constraint (denoted --). Therefore, the joint posterior probability $\hat{{z}}_{{B_{1}},g}^{+\atop-}$ of the pixels in $B_{1}$ belonging to cluster $g$ and the pixels in $B_{2}$ not belonging to cluster $g$ (i.e., given the positive and negative constraints $+\atop-$) is
$\centering\indent\indent\hat{{z}}_{{B_{1}},g}^{+\atop-}=\frac{\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{g}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{g},\hat{\bm{\Lambda}}_{g}\>\hat{\bm{\Lambda}}_{g}^{\top}\>+\>\hat{\bm{\Psi}}_{g})\sum\limits_{\begin{subarray}{c}f=1\\ f\neq g\end{subarray}}^{G}\prod\limits_{\begin{subarray}{c}k\in{B_{2}}\\ k=1\end{subarray}}^{K}\hat{\tau}_{f}\mathcal{N}_{p}(\mathbf{x}_{k};\hat{\bm{\mu}}_{f},\hat{\bm{\Lambda}}_{f}\>\hat{\bm{\Lambda}}_{f}^{\top}\>+\>\hat{\bm{\Psi}}_{f})}{\displaystyle\sum_{g^{{}^{\prime}}=1}^{G}\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{g^{{}^{\prime}}}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{g^{{}^{\prime}}},\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}\>\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}^{\top}\>+\>\hat{\bm{\Psi}}_{g^{{}^{\prime}}})\sum\limits_{\begin{subarray}{c}f=1\\ f\neq g^{{}^{\prime}}\end{subarray}}^{G}\prod\limits_{\begin{subarray}{c}k\in{B_{2}}\\ k=1\end{subarray}}^{K}\hat{\tau}_{f}\mathcal{N}_{p}(\mathbf{x}_{k};\hat{\bm{\mu}}_{f},\hat{\bm{\Lambda}}_{f}\>\hat{\bm{\Lambda}}_{f}^{\top}\>+\>\hat{\bm{\Psi}}_{f})}.\@add@centering$ (4)
Similarly, the joint posterior probability $\hat{{z}}_{{B_{2}},g}^{+\atop-}$ of the pixels in $B_{2}$ belonging to cluster $g$ and the pixels in $B_{1}$ not belonging to cluster $g$ is
$\centering\indent\indent\hat{{z}}_{{B_{2}},g}^{+\atop-}=\frac{\prod\limits_{\begin{subarray}{c}k\in{B_{2}}\\ k=1\end{subarray}}^{K}\hat{\tau}_{g}\mathcal{N}_{p}(\mathbf{x}_{k};\hat{\bm{\mu}}_{g},\hat{\bm{\Lambda}}_{g}\>\hat{\bm{\Lambda}}_{g}^{\top}\>+\>\hat{\bm{\Psi}}_{g})\sum\limits_{\begin{subarray}{c}f=1\\ f\neq g\end{subarray}}^{G}\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{f}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{f},\hat{\bm{\Lambda}}_{f}\>\hat{\bm{\Lambda}}_{f}^{\top}\>+\>\hat{\bm{\Psi}}_{f})}{\displaystyle\sum_{g^{{}^{\prime}}=1}^{G}\prod\limits_{\begin{subarray}{c}k\in{B_{2}}\\ k=1\end{subarray}}^{K}\hat{\tau}_{g^{{}^{\prime}}}\mathcal{N}_{p}(\mathbf{x}_{k};\hat{\bm{\mu}}_{g^{{}^{\prime}}},\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}\>\hat{\bm{\Lambda}}_{g^{{}^{\prime}}}^{\top}\>+\>\hat{\bm{\Psi}}_{g^{{}^{\prime}}})\sum\limits_{\begin{subarray}{c}f=1\\ f\neq g^{{}^{\prime}}\end{subarray}}^{G}\prod\limits_{\begin{subarray}{c}j\in{B_{1}}\\ j=1\end{subarray}}^{J}\hat{\tau}_{f}\mathcal{N}_{p}(\mathbf{x}_{j};\hat{\bm{\mu}}_{f},\hat{\bm{\Lambda}}_{f}\>\hat{\bm{\Lambda}}_{f}^{\top}\>+\>\hat{\bm{\Psi}}_{f})}.\@add@centering$